Let's say the mass of a billiard ball C is 400 g, and the mass of a billiard ball D is 200 g. When ball C struck ball D which was not moving, the velocity of ball C was 6 m/s. If ball C's velocity is reduced to 2 m/s after collision, what is the velocity of ball D after collision? Assume the collision is perfectly elastic and there is no effect of spins, friction, or air resistance.
Answer #1

Dear student,

Find this solution.if any issue with that don't forget to write in comment section.I will rectify them as soon as possible.

If you find the solution helpful and kindly RATE THE ANSWER it would be appreciated.

Your rating is important to me.

Thanks for asking..Mc = 400gmo Mo = 200gm, :ų=6 m V= 2 m using momentum conservation, Meų to = Me Vet Moyo 486x6 50 = 400*2 + 200 XV y = 24-3 =